OpenStudy (anonymous):

a country's population in 1994 was 76 million. In 1998 it was 81 millio. estimate the population in 2016 using the exponential growth formula

OpenStudy (anonymous):

81 - 76 = 5. So it grew 5 million in 4 years. 5,000,000 / 4 = 1250000 1250000 * 18 = 22500000

OpenStudy (vivek3461):

The exponential relation is given by, \[Y = Yo * a^{n}\] Where, Y is the Population at 2016 Yo is the Population at 1994 a is the Growth rate of the population n is the number of years from the starting year to the final year For the above problem n = 12 and Yo = 76000000 Y and a are unknown... To find 'a' use the given data: By substituting the given data for population at 1998, in the exponential relation we get, \[81000000 = 76000000 * a ^{4}\] By solving the above equation you get the value of 'a', With this 'a' using the exponential relation you can find the population at 2016